OpenStudy (anonymous):

suppose you write a book. the printer charges $4 per book to print it and you spend $5500 on advertising. you sell the book for $15 a copy. how many copies must you sell to break even? what is your income for that many (or costs) for that many copies?

OpenStudy (vlery):

Your profit on each book is 15-4=11. 11x=3500 (x being the number of books sold x=318.18, but you can't sell .18 books, so round up to 319

OpenStudy (texaschic101):

5500 + 4b = 15b --- subtract 4b from each side 5500 = 15b - 4b 5500 = 11b 5500/11 = b 500 check.. 5500 + 4(500) = 15(500) 5500 + 2000 = 7500 7500 = 7500 (correct) You would need to sell 500 copies to break even. Costs for that many copies : 5500 + 4(500) = 5500 + 2000 = 7500

OpenStudy (texaschic101):

oh...and income is also $7500